In 12 consecutive innings during a season a batsman’s scores were:

6, 13, 16, 45, 93, 0, 62, 87, 136, 25, 14, 31.

Find his (i) mean score                                                                 


1
Expert's answer
2021-12-07T09:24:39-0500

"\mu=\frac{1}{12}(6+13+16+45+93+0+62+87+136+25+14+31)=\frac{528}{12}=44."